# Foundation Models for the Brain and Body

### v1

@@ A.1 EEGNeX
We used the EEGNeX implementation from the Braindecode package [11, 24]. In this model, we replaced all standard convolutional layers with the convolutional version of our method, Subject-Conditioned Layer (see Appendix 5). We made an exception for the depthwise convolutional layer (DepthwiseConv2D) in Block 3 due to implementation details.
@@
For the PBT, we used the official GitHub repository accompanying the publication and replaced all linear layers with our proposed method, Subject-Conditioned Layer. The only exception was for the visualization in Figure 1, where we kept the final linear layer to improve clarity.
@@
Appendix B Detailed Numerical Results for Subject-Specific Modeling
@@
Appendix C Compute Resources
The shown experiments are performed on a single NVIDIA Tesla V100 GPU using PyTorch. Code is publicly available at github.com/timonkl/SubjectConditionedLayer.

### v2

@@ B.1 Experiment Design
We used the EEGNeX implementation from the Braindecode package (Gramfort et al., 2013; Schirrmeister et al., 2017). In this model, we replaced all standard convolutional layers with the convolutional version of SuLoRA, (see Appendix A). We made an exception for the depthwise convolutional layer (DepthwiseConv2D) in Block 3 due to implementation details.
@@
For the PBT, we used the official GitHub repository accompanying the publication and replaced all linear layers with SuLoRA. The only exception was for the visualization in Figure 2, where we kept the final linear layer to improve clarity.
@@
BrainMagick
For BrainMagick, we used the official GitHub repository accompanying the publication and replaced the subject-specific layer with SuLoRA. We used the same hyperparameters as reported in the original paper for all runs. Specific hyperparameters of our method, respective rank $r$ and $\alpha$ are reported in Table 1.
Table 6: Hyperparameters BrainMagick.
Hyperparameter Value
Signal processing High-pass filter, downsampling
Baseline correction 0.5 s average subtraction
Normalization Robust scaler
Speech representation Wav2vec 2.0 (pretrained)
Subject embedding Yes
Convolutional blocks Stacked CNN architecture
Activation function GELU
Loss function Contrastive (CLIP-style)
Optimizer Adam
Batch size 256
Learning rate $3\times 10^{-4}$
Training epochs Early stopping ($\sim$50% of test set)
Dropout Spatial attention dropout
Segment duration 3 seconds
Appendix C Detailed Numerical Results for Subject-Specific Modeling
@@
Appendix D Software and Data
Upon acceptance, a PyTorch implementation of SuLoRA will be released on GitHub, together with a model wrapper that allows the layer to be easily incorporated into existing architectures, including models from the Braindecode package (Aristimunha et al., 2025b). For the baseline model implementations used in our experiments, we utilized: the EEGNeX implementation from the Braindecode package (Aristimunha et al., 2025b), the PBT implementation from the paper’s GitHub repository (Klein et al., 2025b), and the BrainMagick implementation from the paper’s GitHub repository (Défossez et al., 2023a).
All datasets used in this study are publicly accessible, either directly from OSF (https://osf.io/ag3kj/) or via the BCI Horizon 2020 repository (http://bnci-horizon-2020.eu/database/data-sets). Experiments on BCI datasets are conducted within the MOABB-Package (Chevallier et al., 2024; Aristimunha et al., 2025a).
Appendix E Compute Resources
The experiments for EEGNeX and PBT were performed on a single NVIDIA Tesla V100 GPU using PyTorch. The Brain Magick experiments were performed on single NVIDIA H100 GPU.